OpenStudy (anonymous):

simplify: 5√3(2√3+6√19)

OpenStudy (anonymous):

Use distributive property here..

OpenStudy (anonymous):

Got it :)

OpenStudy (anonymous):

\[\large \color{green}{a \sqrt{b}(c \sqrt{d} + e \sqrt{f}) = a.c \sqrt{b \times d} + a.e \sqrt{b \times f}}\]